Potassium crystallises in a body centred cubic lattice. How many unit cells are present in a 2g potassium?

Grade:12th pass

1 Answers

Umakant biswal
7 years ago
@ vatsala 
no of atoms in the unit cell = 2 ( 1 from body center and 1 from cornor ) 
atomic mass = 39 
no of moles = 2/39 
no of atoms = 2/39 * Na , where ns is avogadro number = 2/39 * 6.022 * 10 ^ 23 
no of unit cell = no of atoms / 2 
= 2* 6.022 * 10 ^ 23 / 2* 39 
so, the no will comes out to be = 1.54 * 10 ^ 22 
which is the answer 
HOPE IT CLEARS YOUR DOUBT 
ALL THE BEST ..